# Geometric Signatures of Compositionality Across a Language Model’s Lifetime

### v2

@@ Abstract
Compositionality, the notion that the meaning of an expression is constructed from the meaning of its parts and syntactic rules, permits the infinite productivity of human language. For the first time, artificial language models (LMs) are able to match human performance in a number of compositional generalization tasks. However, much remains to be understood about the representational mechanisms underlying these abilities. We take a high-level geometric approach to this problem by relating the degree of compositionality in a dataset to the intrinsic dimensionality of its representations under an LM, a measure of feature complexity. We find not only that the degree of dataset compositionality is reflected in representations’ intrinsic dimensionality, but that the relationship between compositionality and geometric complexity arises due to learned linguistic features over training. Finally, our analyses reveal a striking contrast between linear and nonlinear dimensionality, showing that they respectively encode formal and semantic aspects of linguistic composition.
@@ 1 Introduction
By virtue of linguistic compositionality, few syntactic rules and a finite lexicon can generate an unbounded number of sentences (Chomsky, 1957). That is, language, though seemingly high-dimensional, can be explained using relatively few degrees of freedom. A great deal of effort has been made to test whether neural language models (LMs) exhibit human-like compositionality (Hupkes et al., 2019; Baroni, 2019; McCoy, 2022). We take a geometric view of this question, asking how an LM’s representational structure reflects and supports compositional understanding over training.
@@
If a language model is a good model of language, we expect its internal representations to reflect the relatively few variables underlying the latter. That is, representations should reflect the manifold hypothesis, or the notion that real-life, high-dimensional data lie on a low-dimensional manifold (Goodfellow et al., 2016). The dimension of this manifold, or intrinsic dimension (ID), is then the minimal number of degrees of freedom required to describe it without suffering from information loss (Goodfellow et al., 2016; Campadelli et al., 2015). The manifold hypothesis has indeed been attested for linguistic representations: LMs have been found to compress inputs to an ID orders-of-magnitude lower than their extrinsic dimension (Cai et al., 2021; Cheng et al., 2023; Valeriani et al., 2023).
Compositionality permits the atoms of language to locally combine with others, creating global meaning (Frege, 1948; Chomsky, 1999). As such, a complex array of meanings at the level of a phrase is explained by simple rules of composition. A natural question is whether the inherent simplicity of linguistic utterances, enabled by compositionality, manifests in representation manifolds of low complexity, described by the manifolds’ intrinsic dimension. Thus far in the literature, an explicit link between degree of compositionality and representational ID has not been established. To bridge this gap, in a series of controlled experiments on causal language models and a custom dataset with tunable compositionality, we provide the first experimental insights into the relationship between the degree of compositionality of inputs and the ID of their representations over the course of training.
@@
Using our controlledstimulus stimuli and the LMs’ training data, we reproduce the established finding that LMs represent linguistic inputs on low-dimensional, nonlinear manifolds. We also show for the first time that LMs expand representations into high-dimensional linear subspaces, concretely, that (1) nonlinear and linear representational dimension scale differently with model size. We show the relevance of geometry to function over LM training, in particular that (2) LMs’ representational geometry tracks a phase transition in their linguistic competence. Different from past work, we consider two different kinds of compositionality: compositionality of form, or superficial combinatorial complexity, and compositionality of meaning, or semantic complexity; as well as two measures of dimensionality, nonlinear and linear. We not only find that geometric feature complexity reflects input compositionality, but crucially that nonlinear ID encodes meaning compositionality while linear dimensionality encodes form compositionality, in a way that arises over training: (3) nonlinear ID preserves the degree of input compositionality as an inductive bias of the model, but reflects the degree of semantic complexity at the end of training, and (4) linear dimensionality, not nonlinear ID, highly correlates to the superficial combinatorial complexity of inputs. Overall, results reveal a contrast between linear and nonlinear measures of feature complexity that suggests their relevance to form and meaning in how LMs process language.
@@ 2 Background
It has long been a topic of debate whether neural networks also exhibit human-like compositionality when processing natural language (Fodor & Pylyshyn, 1988; Smolensky, 1990; Marcus, 2003). This debate has fueled an extensive line of empirical exploration that assesses the compositionality of neural networks in language modeling via synthetic data (Bentivogli et al., 2016; Lake & Baroni, 2018; Bahdanau et al., 2018) and natural language stimuli (Sathe et al., 2023; Dankers et al., 2022; Press et al., 2023). After the recent introduction of large language models with human-level linguistic capabilities (Wei et al., 2022), researchers have shown via mechanistic interpretability analyses that LMs often extract individual word meanings in early layers, and compose them via later-layer attention heads to construct semantic representations for multi-word expressions (Haviv et al., 2023; Geva et al., 2023). We use complementary tools to understand compositionality: rather than neurons and circuits, we link compositionality to the geometric properties of a model’s embedding space which describe its learned feature complexity.
Language defines a mapping from form to meaning (de Saussure, 1916). Form is the physical shape of an utterance, for example, the sequence of letters or morphemes when written, or sounds when spoken. Broadly, meaning is the concepts or entities to which the forms refer. Unlike prior work, we make a distinction between form and meaning composition, where the formal composition relates to the combinatorial complexity of the data, and semantic composition relates to the ability to construct sentence-level meaning from word meaning. While, in grammatical sentences, meaning composition often inherits from form composition, we disentangle them by creating agrammatical versions of the dataset, further described in the Methods.
The manifold hypothesis and low-dimensional geometry
@@
Deep learning problems are often considered high-dimensional, but research suggests that they have low-dimensional intrinsic structure. In computer vision, studies have shown that common learning objectives and natural image data reside on low-dimensional manifolds (Li et al., 2018; Pope et al., 2021; Valeriani et al., 2023; Psenka et al., 2024). Similarly, learning dynamics of neural LMs have been shown to occur within low-dimensional parameter subspaces (Aghajanyan et al., 2021; Zhang et al., 2023). The nonlinear, low-dimensional structure that emerges in the semantic space of these models, in contrast with models’ tendency to expand representations into high-dimensional linear subspaces (Jazayeri & Ostojic, 2021), has been found to reduce learning complexity (Cheng et al., 2023; Pope et al., 2021), and likely follows from the training objective of predicting sequential observations (Recanatesi et al., 2021).
@@
In the linguistic domain, the geometry of representations has been examined in various contexts. Recent work characterizes the organization of semantic concepts in representation space (Engels et al., 2024; Park et al., 2024; Balestriero et al., 2024; Doimo et al., 2024); it has been found that representational geometry can explicitly encode sparse tree-like syntactic structures (Andreas, 2019; Murty et al., ; Alleman et al., 2021); and that linguistic categories such as part-of-speech are represented in low dimensional linear subspaces (Mamou et al., 2020; Hernandez & Andreas, 2021). Most similar to our setup, Cheng et al. (2023) reported the intrinsic dimension of representations over layers as a measure of feature complexity for several natural language datasets, finding an empirical relationship between information-theoretic and geometric compression. However, our work is the first to explicitly relate the compositionality of inputs, a critical feature of language, to the number of degrees of freedom, or intrinsic dimension, of its representation manifold.
@@
Most research on LMs focuses on the final configuration of the model at the end of pre-training. Yet, recent work shows that learning dynamics can elucidate the behavior and computational mechanisms of LMs (Chen et al., 2024; Singh et al., 2024; Tigges et al., 2024). It has been found that, over training, LMs’ weight matrices become higher-rank (Abbe et al., 2023), their representations higher dimensional (Cheng et al., 2024), and their gradients increasingly diffuse (Weber et al., 2024). Over finetuning, representational dimensionality has been found to change in concert with geometric properties like cluster reorganization (Doimo et al., 2024).
Phase transitions during LM training have been found for some, but not all, aspects of language learning. Negative evidence includes that LM circuits involved in linguistic subtasks are stable (Tigges et al., 2024) and gradually reinforced (Weber et al., 2024) over training. Positive evidence for learning phase transitions includes that the ID of BERT’s final [CLS] representation tracks sudden syntax acquisition and drops in training loss (Chen et al., 2024), with similar observations on Transformers trained on formal languages (Lubana et al., 2024). Our work supplements these results by investigating how the interaction between compositional understanding of language and geometric complexity of its representation arises over training.
@@ 3 Setup
We consider the relationship between a dataset’s degree of compositionality and its representational complexity under an LM. Here, we describe the models, dataset generation, compositionality quantification, and feature complexity estimation.
@@ 3.1 Models
For the three intermediate sizes 410m, 1.4b, and 6.9b, we report model performance throughout the pre-training phase on the set of evaluation suites provided by (Biderman et al., 2023; Gao et al., 2024), further described in Appendix F. This encompasses a range of higher-level linguistic and reasoning tasks, spanning from long-range text comprehension (Paperno et al., 2016) to commonsense reasoning (Bisk et al., 2019). The evolution of task performance provides a cue for the type of linguistic knowledge learned by the model by a certain training checkpoint.
@@ 3.2 Datasets
As we consider the relationship between the degree of compositionality and geometric feature complexity, we create a custom grammar whose compositional structure we can control. In addition, we replicate experiments on The Pile in order to compare results to a general slice of natural language.
@@ 3.2.1 Controlled grammar
Our stimulus dataset consists of grammatical nonce sentences from the grammar illustrated in Figure 1. To create the grammar, we set 12 semantic categories and randomly sample a vocabulary of 50 words for each category, where the categories’ vocabularies are disjoint. The categories include 5 adjective types (quality, nationality, size, color, texture), 2 noun types (job, animal) and 1 verb type. We use a simple, fixed syntactic structure by ordering the word categories:
@@
This produces sentences that are $16$ words long. The order is chosen so that the generated sentences are grammatical and that the adjective order complies with the accepted ordering for English (Dixon, 1976). Vocabularies are chosen such that the sentences are semantically coherent. For example, for the first verb, the agent is a person and patient is an animal, so the possible verbs are constrained to permit “walks", but not “types". The vocabularies for each category may be found in Appendix E.
Figure 1: Dataset structure and distributional properties. Top: The structure of the stimulus dataset. The top row shows the ordering of word categories, such as quality.ADJ or animal.N; below it, the vocabulary for each category, including words like “strong" (quality.ADJ) and “lion" (animal.N), respectively. When controlling the degree of dataset compositionality, contiguous word positions are coupled. For instance, when $k=3$, the first vocabulary indices for quality1.ADJ, nationality1.ADJ, and job1.N are tied together, such that “strong Thai doctor" or “calm French teacher" can be sampled, but “strong French doctor" cannot. Left: Examples of generated prompts for the normal, $k=3$, and shuffled settings. Right: When controlling the compositionality across $k=1\cdots 4$, word unigram frequencies are preserved in the resulting datasets, shown in the distributions looking identical.
Although the syntactic structure and individual vocabulary items are likely seen during training, words are sampled independently for each category without considering their probability in relationship to other words in the sentence. Therefore, generated sentences are highly unlikely to be in the training data. [^1] We cannot verify that utterances aren’t in the training set, as at the time of submission, it is not possible to search The Pile. Then, when encountering these sentences for the first time, a frozen LM must successfully construct their meanings from the meanings of their parts, or compositionally generalize.
@@
We modify the grammar in order to vary the dataset’s degree of compositionality. While linguistic compositionality spans many interpretations (Hupkes et al., 2019), [^2] We do not consider the recursive, hierarchical nature of compositionality theorized by Chomskian linguists. We leave, e.g., different levels of syntactic embedding to future work. we are interested in two specific types: (1) composition of forms, or combinatorial complexity of the dataset, where a dataset is more compositional if it contains more unique word combinations; (2) composition of meanings, or sentence-level compositional semantics, where sentence meaning is composed, via syntax, from word meanings.
First, to control for dataset combinatorial complexity, we couple the values of $k$ contiguous word positions for $k=1\cdots 4$. That is, the sequence’s atomic units are sets of $k$ adjacent words, or $k$-grams, sampled independently. This constrains the degrees of freedom in sampling to $l/k$ where $l=12$ is the number of categories: for instance, in the $1$-coupled setting, each word is sampled independently, hence 12 degrees of freedom; in the $2$-coupled setting, each bigram is sampled independently, hence 6 degrees of freedom. Varying $k$ maintains the dataset’s unigram distribution by design (see Figure 1 right), but constrains the dataset’s $k$-gram distributions, or combinatorial complexity.
To investigate compositional semantics, we randomly shuffle the words in each sequence. This destroys syntactic coherence, and in turn, the overall meaning of the sentence. It instead preserves superficial distributional properties like word count and word co-occurrences at the sentence level, as well as unigram frequencies (see Figure 1 right). Then, LM behavior on grammatically sane vs. shuffled sequences proxies compositional vs. lexical-only semantics.
@@
Measuring formal and semantic compositionality
@@
Meaning complexity differs from form complexity. For example, the data [cat, lion, puma] are related semantically but not formally. As there is no unified definition for semantic complexity (Pollard & Biermann, 2000; Chersoni et al., 2016), we do not attempt to quantify it. But, as sane sequences are grammatical and semantically coherent, it is guaranteed for sane datasets that meaning complexity is monotonic in form complexity. In addition, as shuffling removes sequence-level semantics, meaning complexity is guaranteed to be lower on shuffled compared to sane text, by definition.
@@ 3.2.2 The Pile
Although we focus on the controlled grammar in order to vary compositionality, to ensure that results are not an artifact of our prompts, we replicate experiments on The Pile, a general slice of natural language consisting of encyclopedic text, social media, reviews, news articles, and books. We uniformly sample $N=50000$ sequences in The Pile, each consisting of $16$ words, the same length as sequences in the controlled grammar, and report results over $5$ random data splits.
3.3 Measuring feature complexity via dimensionality estimation
We are interested in how the geometric complexity of representations reflects the inputs’ degree of compositionality. In particular, we consider representations in the Transformer’s residual stream (Elhage et al., 2021). Because sequence lengths may slightly vary due to the tokenization scheme, in line with prior work (Cheng et al., 2023; Doimo et al., 2024), we aggregate over the sequence by taking the last token representation, as, due to causal attention, it is the only to attend to the entire context.
For each layer and dataset, we compute both a nonlinear and a linear measure of dimensionality. Nonlinear and linear dimensionality have key conceptual differences. The nonlinear $I_{d}$ is the number of degrees of freedom, or latent features, needed to describe the underlying manifold (Campadelli et al., 2015; Facco et al., 2017), see Appendix D for discussion. This differs from the linear effective dimensionality $d$, the dimension of the minimal linear subspace that contains the set of representations. Throughout, we will use dimensionality to refer to both nonlinear and linear estimates. When appropriate, we will specify $I_{d}$ as the nonlinear ID, $d$ as the linear effective dimensionality, and $D$ as the extrinsic dimensionality, or hidden dimension of the model. Since an $I_{d}$-dimensional manifold can be embedded in a $\geq I_{d}$-dimensional linear subspace, we always have that $I_{d}\leq d\leq D$.

### v3

@@ Abstract
By virtue of linguistic compositionality, few syntactic rules and a finite lexicon can generate an unbounded number of sentences. That is, language, though seemingly high-dimensional, can be explained using relatively few degrees of freedom. An open question is whether contemporary language models (LMs) reflect the intrinsic simplicity of language that is enabled by compositionality. We take a geometric view of this problem by relating the degree of compositionality in a dataset to the intrinsic dimension (ID) of its representations under an LM, a measure of feature complexity. We find not only that the degree of dataset compositionality is reflected in representations’ ID, but that the relationship between compositionality and geometric complexity arises due to learned linguistic features over training. Finally, our analyses reveal a striking contrast between nonlinear and linear dimensionality, showing they respectively encode semantic and superficial aspects of linguistic composition.
Geometric Signatures of Compositionality
Across a Language Model’s Lifetime
Jin Hwa Lee1, Thomas Jiralerspong2,3, Lei Yu4, Yoshua Bengio2,3, Emily Cheng5 1University College London, 2Université de Montréal, 3Mila, 4University of Toronto, 5Universitat Pompeu Fabra Correspondence: jin.lee.22@ucl.ac.uk and thomas.jiralerspong@mila.quebec and emilyshana.cheng@upf.edu
@@ 1 Introduction
Compositionality, the notion that the meaning of an expression is constructed from that of its parts and syntactic rules (Szabó, 2024), enables linguistic atoms to locally combine to create global meaning (Frege, 1948; Chomsky, 1999). Then, a rich array of meanings at the level of a phrase may be explained by simple rules of composition. If a language model is a good model of language, we expect its internal representations to reflect the latter’s relative simplicity. That is, representations should reflect the manifold hypothesis, or that real-life, high-dimensional data lie on a low-dimensional manifold (Goodfellow et al., 2016).
@@
The dimension of this manifold, or intrinsic dimension ($I_{d}$), is the minimal number of degrees of freedom needed to describe it without information loss (Goodfellow et al., 2016; Campadelli et al., 2015). The manifold hypothesis is attested for linguistic representations: LMs are found to compress inputs to an $I_{d}$ orders-of-magnitude lower than their ambient dimension (Cai et al., 2021; Cheng et al., 2023; Valeriani et al., 2023).
A natural question is whether the inherent simplicity of linguistic utterances, enabled by compositionality, manifests in representation manifolds of low complexity, described by the manifolds’ $I_{d}$. For example, consider the set of possible phrases generated by the syntactic structure [DET][N][V]. While the possible utterances have combinatorial complexity, they are generated by only three degrees of freedom. An LM that can handle this syntactic structure should be able to both (1) represent all combinations; (2) encode the few degrees of freedom underlying them, or their degree of compositionality. We ask, which geometric aspects of the representation space encode (1) and (2)?
The manifold hypothesis predicts the $I_{d}$ of LM representations to encode (2) (Recanatesi et al., 2021). But, so far in the literature, while some linguistic categories have been found to occupy low-dimensional linear subspaces (Hernandez and Andreas, 2021; Mamou et al., 2020), an explicit link between degree of linguistic compositionality and representational dimensionality has not been established. To bridge this gap, in large-scale experiments on causal LMs and a custom dataset with tunable compositionality, we provide first insights into the relationship between the degree of compositionality of inputs and geometric complexity of their representations over the course of training.
Outline
@@
On both our controlled datasets and naturalistic stimuli, we reproduce the finding (Cai et al., 2021) that LMs represent linguistic inputs on low-dimensional nonlinear manifolds. We also show for the first time that LMs expand representations into high-dimensional linear subspaces, concretely, that (I) nonlinear and linear dimensionality scale differently with model size. We show that feature geometry is relevant to behavior over training, in that (II) LMs’ feature complexity tracks a phase transition in linguistic competence. Different from past work, we consider two different measures of dimensionality, nonlinear and linear, showing that they encode complementary information about compositionality. Our key result is that representations’ geometric complexity reflects dataset complexity. However, nonlinear $I_{d}$ encodes (2) meaningful compositionality while linear dimensionality encodes (1) superficial input complexity, in a way that arises over training: (III) nonlinear $I_{d}$ reflects superficial complexity as an inductive bias of the newly initialized LM, but meaningful compositional complexity by the end of training. Instead, (IV) linear dimensionality, not $I_{d}$, encodes superficial complexity throughout training. Overall, results show a contrast between linear and nonlinear feature complexity that suggests their relevance to form and meaning in how LMs process language.
@@ 2 Background
Whether neural networks exhibit human-like compositionality when processing natural language is an open question (Fodor and Pylyshyn, 1988; Smolensky, 1990; Marcus, 2003). This question has fueled extensive research that assesses neural networks’ compositionality in language modeling via synthetic data (Bentivogli et al., 2016; Lake and Baroni, 2018; Bahdanau et al., 2018) and natural language stimuli (Sathe et al., 2023; Dankers et al., 2022; Press et al., 2023). After the recent introduction of LMs with human-level linguistic capabilities (Wei et al., 2022), researchers have shown, via mechanistic interpretability, that LMs extract word meanings in early layers and compose them in later ones to construct the meaning of multi-word expressions (Haviv et al., 2023; Geva et al., 2023). We use complementary tools to understand compositionality: instead of neurons and circuits, and rather than explicitly describe the LM’s composition function (Smolensky, 1990), we are interested in the extent to which LMs preserve inputs’ compositional complexity in their representations’ geometric complexity.
Manifold hypothesis and low-dimensional geometry
@@
Deep learning problems appear high-dimensional, but research suggests that they have low-dimensional intrinsic structure. In computer vision, image data and common learning objectives live on low-dimensional manifolds (Li et al., 2018; Pope et al., 2021; Valeriani et al., 2023; Psenka et al., 2024; Ansuini et al., 2019). Similarly, LMs’ learning dynamics occur in low-dimensional parameter subspaces (Aghajanyan et al., 2021; Zhang et al., 2023). The parsimony that governs these models’ latent spaces reduces learning complexity (Cheng et al., 2023; Pope et al., 2021) and may arise from the training objective, seen for classification (Chung et al., 2018) and sequential prediction (Recanatesi et al., 2021).
@@
For language, the geometry of representations has been explored in several contexts. Prior work describes how concepts organize in representation space (Engels et al., 2024; Park et al., 2024; Balestriero et al., 2024; Doimo et al., 2024); it has been found that representational geometry encodes tree-like syntactic structures (Andreas, 2019; Murty et al., ; Alleman et al., 2021; Hewitt and Manning, 2019); and that linguistic categories like part-of-speech lie in low-dimensional linear subspaces (Mamou et al., 2020; Hernandez and Andreas, 2021). Most similar to our work, Cheng et al. (2023) report the $I_{d}$ of representations over layers as a measure of feature complexity for natural language datasets, finding an empirical link between information-theoretic and geometric compression. But, our work is the first to explicitly relate the compositionality of inputs, a critical feature of language, to the degrees of freedom, or $I_{d}$, of their representation manifold.
@@
Most research on LMs focuses on their final configuration at the end of pre-training. Yet, recent work shows that learning dynamics provide cues as to how behavior arises (Chen et al., 2024; Singh et al., 2024; Tigges et al., 2024). It has been found that, over training, LM weights become higher-rank (Abbe et al., 2023), their features higher-dimensional (Cheng et al., 2025), and their gradients more diffuse (Weber et al., 2024). Over finetuning, feature complexity changes in concert with geometric properties like clustering (Doimo et al., 2024).
Phase transitions in LM training are attested for some, but not all, aspects of language learning. Negative evidence includes that circuits involved in linguistic tasks are stable (Tigges et al., 2024) and gradually reinforced (Weber et al., 2024) over training. Positive evidence includes that BERT´s feature complexity tracks sudden syntax acquisition and drops in training loss (Chen et al., 2024), with similar findings for Transformers trained on formal languages (Lubana et al., 2024). Our work complements these results, exploring how the relationship between compositionality of inputs and complexity of their representations arises over training.
@@ 3 Setup
We consider the relationship between a dataset’s degree of compositionality and its feature complexity under an LM. Here, we describe the models and datasets, as well as how to measure dataset compositionality and feature complexity.
@@ 3.1 Models
Intermediate training checkpoints are public for Pythia models (Biderman et al., 2023). For three intermediate Pythia sizes 410m, 1.4b, and 6.9b, we report model performance throughout the pre-training phase on the set of evaluation suites provided by (Biderman et al., 2023; Gao et al., 2024), detailed in Appendix F. This encompasses a range of higher-level linguistic and reasoning tasks, spanning from long-range text comprehension (Paperno et al., 2016) to commonsense reasoning (Bisk et al., 2019). The evolution of task performance provides a cue for the type of linguistic knowledge learned by the model by a certain training checkpoint.
@@ 3.2 Datasets
As we consider the relationship between the degree of compositionality and geometric feature complexity, we create a custom grammar whose compositional structure we can control. We additionally replicate key experiments on The Pile so to compare results to a general slice of natural language.
@@ 3.2.1 Controlled grammar
Our stimulus dataset consists of grammatical sentences generated from a custom grammar. To create it, we set $12$ semantic categories and uniformly sample a $50$-word vocabulary for each category; different categories’ vocabularies are disjoint. The categories include $5$ adjective types (quality, nationality, size, color, texture), $2$ noun types (job, animal) and $1$ verb type. We use a fixed syntactic structure by ordering the word categories:
@@
This produces natural language sentences that are $17$ words long. The order is chosen so that generated sentences are grammatical and that the adjective order complies with the accepted ordering for English (Dixon, 1976). Vocabularies are chosen so that sentences are semantically coherent: for instance, if the agent is a person and patient is an animal, then verbs are constrained to permit, e.g., “pets", but not “types". Further, to explore the effect of sequence length on feature complexity, we design similar grammars producing sentences of other lengths $\in\{5,8,11,15\}$. The detailed sampling procedure, vocabularies, and structures of the different grammars are found in Appendix E.
Although the syntactic structure and individual vocabulary items are likely seen during training, words are sampled independently for each category without considering their probability in relationship to other words in the sentence. Therefore, generated sentences are highly unlikely to be in the training data, thus memorized by the LM. [^1] We can’t verify utterances aren’t in the training set, as at the time of submission, it is not possible to search The Pile. Then, when processing these sentences, a frozen LM must successfully construct their meanings from the meanings of their parts, or compositionally generalize.
@@
We modify the grammar to vary the dataset’s degree of compositionality. To do so, we couple the values of $k$ contiguous word positions to produce a different dataset for each $k=1\cdots 4$. That is, during data generation, $k$-grams are independently sampled, which constrains the degrees of freedom in each sentence to $l/k$ where $l$ is the number of categories. For instance, for the longest grammar, there are $l$=$12$ categories; in the $1$-coupled setting, each category is sampled independently, hence $12$ degrees of freedom; in the $2$-coupled setting, each bigram is sampled independently, hence $6$ degrees of freedom. Varying $k$ maintains the dataset’s unigram distribution, but constrains the dataset’s $k$-gram distributions, or compositional complexity.
Beyond varying $k$, we perform a severe intervention to inputs’ compositional structure where we randomly shuffle the words in each sequence. This breaks inputs’ syntactic coherence, and in turn, the overall meaning of the sentence. It rather preserves superficial distributional properties like length, co-occurrences at the sentence level, and unigram frequencies (Figure E.1 right). Then, LM processing of grammatically coherent vs. shuffled text proxies compositional vs. bag-of-words semantics.
@@
Measuring compositionality
@@
The connection between a dataset’s KC and its degree of compositionality requires careful interpretation, as gzip is not semantics-aware: gzip only measures the compressibility of the superficial wordforms that make up the sequences. In coherent datasets, sequences are grammatical and literal (i.e., semantic composition of wordforms is regular and predictable); then, the superficial complexity measure returned by gzip directly proxies the degree of compositionality. For shuffled datasets, whose sequences do not permit a compositional interpretation, the KC instead only characterizes wordforms’ superficial combinatorial complexity.
@@ 3.2.2 The Pile
We focus on the synthetic grammar to tune compositionality in a surgical way. But, to ensure results are not an artifact of our prompts, we replicate experiments on The Pile, uniformly sampling $5$ random splits of $N$=$10000$ sequences that are $17$ words, the same length as sequences in the longest controlled grammar. We report results over splits.
3.3 Measuring feature complexity
We are interested in how feature complexity reflects the inputs’ degree of compositionality. In particular, we consider representations in the Transformer’s residual stream (Elhage et al., 2021). As sequence lengths can slightly vary due to tokenization, as in prior work (Cheng et al., 2023; Doimo et al., 2024), we aggregate over the sequence by taking the last token representation, as, due to causal attention, it is the only one to attend to the entire context.
For each layer and dataset, we measure feature complexity using both a nonlinear and a linear measure of dimensionality. Nonlinear and linear measures have key differences. The nonlinear $I_{d}$ is the number of degrees of freedom, or latent features, needed to describe the underlying manifold (Campadelli et al., 2015) (see Appendix C). This differs from the linear effective dimensionality $d$, the dimension of the minimal linear subspace that explains representations’ variance. Throughout, we will use dimensionality to refer to both nonlinear and linear estimates. When appropriate, we will specify $I_{d}$ as the nonlinear ID, $d$ as the linear effective dimension, and $D$ as the ambient dimension, or the LM’s hidden dimension. An $I_{d}$-dimensional manifold can be embedded in $\mathbb{R}^{d}$ where $d\geq I_{d}$ (Munkres, 2000), hence $I_{d}\leq d\leq D$ by definition.
[…]
Figure 1: Dimensionality over model size. Mean layerwise $I_{d}$ (left) and $d$ (right) are shown for increasing hidden dimension. $I_{d}$ does not depend on hidden dimension $D$ (flat lines), but PCA $d$ scales linearly in $D$. Curves are averaged over 5 data splits $\pm$ 1 SD.